The following equation gives the amount of money owed on a loan after a certain amount of time if no payments are made. A = P ( 1 + r/n ) ^nt where:

A = the amortized amount (total loan/investment amount over the life of the loan/investment)

P = the initial amount of the loan/investment

r = the annual rate of interes

t n = the number of times interest is compounded each year

t = the time in years

Find the amount owed at the end of 2 years if $ 3 , 500 is loaned at a rate 7 % compounded monthly. round answer 2 decimal places
Answer #1

Sol: Griven P. = $3500, t= 2 4 cons N-70% = 0.07 - 12. Fermula: 7. Р (Hoht 1272 =-3500(1+ ono 0.07 $21024.32076. A ~ $14024.3